OpenStudy (anonymous):

x^2 + 2x=3.....i have totally forgotten how to solve things like this. its the exponents that are throwing me off. help? :(

OpenStudy (anonymous):

well you can either factor the quadratic equation or you can use a formulea.... which one do you want me to explain?

OpenStudy (anonymous):

ummm unfortunately i cant remember how to do either :(

OpenStudy (anonymous):

hahaha okay well personally i fine the formulea method easiest because all you have to do is substitute in the variables. the formulea is x=\[(-b \pm \sqrt{(b)^2-4(a)(c)}) /2(a)\] . so your equation is x^2+2x=3. first you take the 3 over to the other side of the = sign and the equation therefore becomes x^2+2x-3=0. so we will name the coefficient of x^2 as a, the coefficient of 2x as b and the coefficient of -3 as c. ( a coefficient is the number in front of the x. therefore x^2 coefficient is 1 which =a, coefficient of 2x =2=b and the coefficient of -3=-3=c) so... all you do is take your coefficients and put them into that equation ( you can type that equation into you calculator but dont forget the brackets) and u will get your answer. BUT REMEMBER... when you type that into your calc, you have to change the +- sign into a +, get an answer then type the whole thing out again and replace the +- sign with a - the second time around and get an answer. hence, you will have 2 answers

OpenStudy (anonymous):

your answers should be x=-3 and x=1

OpenStudy (anonymous):

oh my gosh thank you so much. i'm kind of starting to remember it now. I never did well with factoring. THANK YOU SOOOOO MUCH

OpenStudy (anonymous):

yeah i hate factoring too... :) its a pleasure